OpenStudy (anonymous):

Help this is super confusing!! Write the recursive formula for the geometric sequence. ^a1=-2 ^a2=8 ^a3=-32

hartnn (hartnn):

did you notice, a2=a1*(-4) a3=a2*(-4) ??

hartnn (hartnn):

so the recursive formula u can make is \[a _{n}=a _{n-1}*-4\]